A canoe trip upstream to a campground takes six hours.
The trip back downstream takes three hours.
The current flows at 3 miles per hour.
Write an equation for total roundtrip distance travelled (d) in terms of your still water speed.

Answer by richwmiller(17219)

same distance
r*t=d
(r-3)*6=(r+3)*3
6r-18=3r+9
3r=27
r=9
2*r*t=d
2*(9+3)*3=d
2*12*3=d
72 is round trip distance